# Mysterious Projections: Multimodal LLMs Gain Domain-Specific Visual Capabilities Without Richer Cross-Modal Projections

### v1

@@ 1 Introduction
In this work, we use domain-specific fine-tuning using the above two strategies to understand the role of the projection and the LLM parameters in acquiring domain-specific image modeling capabilities. We posit that if the projection plays a critical role in acquiring domain-specific image modeling capabilities, the post-projection representation – i.e., the representation of the image transformed by the projection, should be richer [^2] We use domain-specific richness to indicate the “expressive power” of the representations Bengio et al. (2012) towards the domain-specific task. in domain-specific features. Conversely, if the post-projection representation is not richer in domain-specific features, the domain-specific features are being identified or modeled by the LLM parameters. [^3] Project webpage: https://claws-lab.github.io/projection-in-MLLMs
@@ 3 Role of Projection in Learning Domain-Specific Image Attributes
Comparing domain-specific richness of post-projection representation across different settings. Table 2 shows: (a) the domain-specific richness of post-projection representation $\mathbf{H_{v}}$ (‘Post-proj MLP’), and (b) the corresponding MLLM performance (‘MLLM’), across the three settings mentioned above (i.e., ‘Original’, ‘FT-Proj’, and ‘FT-E2E’). We report the macro-averaged $F_{1}$ score on the test set of the respective dataset for both (a) and (b). There are two key trends in Tablre 2: first, when the ‘Original’ LLaVA-1.5 model’s projection layer is fine-tuned (‘FT-Proj’), the domain-specific richness of the post-projection representation diminishes, while a boost in the MLLM performance is observed. Similarly, second, with end-to-end fine-tuning of LLaVA-1.5 (‘FT-E2E’), the domain-specific richness of the post-projection representation worsens while the MLLM performance boosts notably. These two trends are consistent across all the datasets considered in our study.
@@ 5 Limitations and Broader Perspective
Datasets and code: The datasets used in this study are publicly available and were curated by previous research. The code an resources associated with the work are available at: https://claws-lab.github.io/projection-in-MLLMs.
@@ A.2 Multilayer Perceptron for Estimating Post-Projection Richness
We train a multilayer perceptron for estimating the domain-specific richness of the post-projection image representation (i.e., $\textbf{H}_{v}$). The MLP takes the 4096 tokens corresponding to the image as input and learns to perform the classification task using the examples from the standard train set. Architecturally, the MLP comprises a token-level average pooling step to obtain 4096-dimensional input vector, followed by subsequent layers of size $2048$, $1536$, $1024$, $512$, $256$, and the output layer of size equivalent to the number of classes in the dataset. We use ReLU activation Agarap (2018) to induce non-linearity. We keep the architecture of this MLP fixed across all the settings to control for the number of learnable parameters and the representational power of the neural network, therefore allowing us to estimate the richness of the input embeddings with respect to the target task. Each model is trained with a batch size of 128. We use Adam optimizer  Kingma and Ba (2014) with a learning rate initialized at $10^{-4}$ and adopt early stopping based on the loss values to avoid overfitting. As a sanity check, we note that an MLP trained using our setup on the post-projection embeddings obtained from the original LLaVA-1.5 model for the Humanitarian task (a natural images dataset), achieves close to the state-of-the-art performance reported on this task Alam et al. (2018). This indicates that our setup enables a reliable estimate of the richness/expressive power of the post-projection representations.

### v2

# Cross-Modal Projection in Multimodal LLMs Doesn’t Really Project Visual Attributes to Textual Space

@@ 1 Introduction
In this work, we use domain-specific fine-tuning using the above two strategies to understand the role of the projection and the LLM parameters in acquiring domain-specific image modeling capabilities. We posit that if the projection plays a critical role in acquiring domain-specific image modeling capabilities, the post-projection representation – i.e., the representation of the image transformed by the projection, should be richer [^2] We use domain-specific richness to indicate the “expressive power” of the representations Bengio et al. (2012) towards the domain-specific task. in domain-specific features. Conversely, if the post-projection representation is not richer in domain-specific features, the domain-specific features are being identified or modeled by the LLM parameters. [^3] Project webpage: https://claws-lab.github.io/projection-in-MLLMs/
@@ 3 Role of Projection in Learning Domain-Specific Image Attributes
Comparing domain-specific richness of post-projection representation across different settings. Table 2 shows: (a) the domain-specific richness of post-projection representation $\mathbf{H_{v}}$ (‘Post-proj MLP’), and (b) the corresponding MLLM performance (‘MLLM’), across the three settings mentioned above (i.e., ‘Original’, ‘FT-Proj’, and ‘FT-E2E’). We report the macro-averaged $F_{1}$ score on the test set of the respective dataset for both (a) and (b). There are two key trends in Table 2: first, when the ‘Original’ LLaVA-1.5 model’s projection layer is fine-tuned (‘FT-Proj’), the domain-specific richness of the post-projection representation diminishes, while a boost in the MLLM performance is observed. Similarly, second, with end-to-end fine-tuning of LLaVA-1.5 (‘FT-E2E’), the domain-specific richness of the post-projection representation worsens while the MLLM performance boosts notably. These two trends are consistent across all the datasets considered in our study.
@@ 5 Limitations and Broader Perspective
Datasets and code: The datasets used in this study are publicly available and were curated by previous research. We abide by their terms of use. We release the code for our experiments to aid reproducibility and enable future research on this topic: https://github.com/claws-lab/projection-in-MLLMs
6 Acknowledgements
This research/material is based upon work supported in part by NSF grants CNS-2154118, ITE-2137724, ITE-2230692, CNS2239879, Defense Advanced Research Projects Agency (DARPA) under Agreement No. HR00112290102 (subcontract No. PO70745), CDC, and funding from Microsoft. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the position or policy of DARPA, DoD, SRI International, CDC, NSF, and no official endorsement should be inferred. Gaurav is partially supported by the JP Morgan AI Research PhD Fellowship and the Snap Research Fellowship. We thank the members of the CLAWS Lab for their helpful feedback.
@@ A.2 Multilayer Perceptron for Estimating Post-Projection Richness
We train a multilayer perceptron for estimating the domain-specific richness of the post-projection image representation (i.e., $\textbf{H}_{v}$). The MLP takes the tokens corresponding to the image as input and learns to perform the classification task using the examples from the standard train set. Architecturally, the MLP comprises a token-level average pooling step to obtain the image representation, followed by subsequent layers, and eventually the output layer of size equivalent to the number of classes in the dataset. We use ReLU activation Agarap (2018) to induce non-linearity. We keep the architecture of this MLP fixed across all the settings to control for the number of learnable parameters and the representational power of the neural network, therefore allowing us to estimate the richness of the input embeddings with respect to the target task. Each model is trained with a batch size of 128. We use Adam optimizer  Kingma and Ba (2014) with a learning rate initialized at $10^{-4}$ and adopt early stopping based on the loss values to avoid overfitting. As a sanity check, we note that an MLP trained using our setup on the post-projection embeddings obtained from the original LLaVA-1.5 model for the Humanitarian task (a natural images dataset), achieves close to the state-of-the-art performance reported on this task Alam et al. (2018). This indicates that our setup enables a reliable estimate of the richness/expressive power of the post-projection representations.